# LEGO_HQEC: A Software Tool for Analyzing Holographic Quantum Codes

### v1

@@ Abstract
Quantum error correction (QEC) is a crucial prerequisite for future large-scale quantum computation. Finding and analyzing new QEC codes, along with efficient decoding and fault-tolerance protocols, is central to this effort. Holographic codes are a recent class of QEC subsystem codes derived from holographic bulk/boundary dualities. In addition to exploring the physics of such dualities, these codes possess useful QEC properties such as tunable encoding rates, distance scaling competitive with topological codes, and excellent recovery thresholds. To allow for a comprehensive analysis of holographic code constructions, we introduce LEGO_HQEC, a software package utilizing the quantum LEGO formalism. This package constructs holographic codes on regular hyperbolic tilings and spawns their stabilizer generators and logical operators for a specified number of seed codes and layers. Three decoders are included: an erasure decoder based on Gaussian elimination; an integer-optimization decoder; and a tensor-network decoder. With these tools, LEGO_HQEC thus enables future systematic studies regarding the utility of holographic codes for practical quantum computing.
@@ I Introduction
The field of quantum computing is currently focused on understanding how best to construct a large-scale fault-tolerant quantum computer [1, 2, 3, 4, 5, 6]. The physical levels of noise that dominate in current quantum computing hardware are simply too high to reliably run all but the simplest quantum algorithms. As a solution quantum error correction (QEC) techniques [7, 8, 9, 10] were proposed in order to orchestrate strategies that protect information at the logical level and facilitate the execution of a quantum algorithm on a device.
@@
Although QEC in theory posits a method by which practically useful quantum computing could be realized, in reality the constraints imposed by current state-of-the-art code constructions and so-called fault-tolerance protocols are extremely resource-intensive [4]. As such, almost all areas of the QEC design pipeline are under active research. Some of these research areas range from the theoretical to the practical, and include: code construction [11, 12, 13]; fault tolerant error correction protocols [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]; fault-tolerant universal logical gate set implementations, such as magic-state distillation [26], gauge-fixing [27], and code switching [28, 29, 30], among others [31, 32]; decoder design [33, 34, 35, 36]; architectural proposals [1, 37, 38]; among many other engineering-level challenges [39, 12].
@@
Holographic quantum codes are a recently introduced class of quantum error correction codes [40, 41, 13], derived as models of bulk/boundary dualities in theoretical physics, in particular the AdS/CFT correspondences [42, 43]. The main idea for these codes lies in the fact that bulk degrees of freedom, which are associated with logical qubits in $d$ dimensions, are mapped to a $(d-1)$-dimensional boundary, with the hyperbolic structure of the bulk ensuring that the physical Hilbert space is larger than its logical counterpart. Constructed on hyperbolic manifold tessellations, they can be expressed as concatenated subsystem codes and allow for extremely high (and tunable) encoding rates as well as excellent threshold performance in comparison to other leading quantum code constructions, such as quantum low-density parity-check (qLDPC) [12] and topological (TQEC) quantum codes [39, 37]. However, for all practical purposes, actually generating the conformal quasicrystal lattices [44, 45, 46] responsible for defining physical-qubit representations of stabilizers and logical operators is a challenging task, exacerbated by the need to apply suitable decoding algorithms derived from tensor-network decoding techniques [47, 48, 49, 50]. In this vein, having a software tool at hand to compute the effect of quantum noise channels is indispensable for the study of new holographic (as well as generalized concatenated) quantum codes.
@@
In this work, we introduce a software tool named LEGO_HQEC, which has been designed to ease the code- and operator-generation process, thereby expediting and lowering the barrier to entry for this new subfield of quantum error correction. In particular, we implement a stabilizer and logical operator generator that returns a list of said objects for both maximum-rate and asymptotically zero-rate variants of many existing holographic codes, as well as providing valuable customization options for constructing homogeneous or heterogeneous constructions (i.e., holographic codes constructed from either one or several smaller seed codes) on hyperbolic manifolds parameterized by Schläfli symbols $\{p,q=4\}$ [51]. In addition to this tool, we provide three different decoders which have been used in recent studies regarding the threshold properties of holographic codes for practical quantum computing: an erasure decoder built using a Gaussian elimination (GE) method [52, 53]; a Pauli-noise integer-optimization decoder, as from [54, 52]; and a tensor-network decoder, which was utilized in [55, 48, 49] additionally for Pauli noise. With the aim of developing newer and better open-source software tools for studying quantum error correction, this package has been made publicly available at https://github.com/QML-Group/HQEC.
@@
This paper is organized as follows: Section II gives an overview of the quantum LEGO formalism, and how it relates to stabilizer and holographic quantum error correction codes; Section III details the structure of LEGO_HQEC, starting with boundary stabilizer and logical operator generation, and ending with a concise overview of the distinct decoders offered; in Section IV, we show two examples using: the $\llbracket 5,1,3\rrbracket$ (or Harlow-Preskill-Pastawski-Yoshida (HaPPY) [40]), as well as a holographic Steane code [53]. The calculated runtime for generating stabilizers and logical operators for several layers of the asymptotically zero-rate HaPPY code is presented, as well as decoding for erasure and Pauli noise in the asymptotically zero-rate Steane code, in addition to the maximum-rate HaPPY code. We conclude by proposing several extensions to LEGO_HQEC in Section V.
@@ II Background
II.1 Quantum Error Correction & Tensor Networks
A quantum error correction (QEC) code is a technique used to protect information at the level of physical qubits (or qudits) from deleterious environmental noise effects. Such protection is accomplished by redundantly mapping physical qubit states into subspaces of a larger, multipartite Hilbert space; this description at the level of an encoding isometry effectively defines a QEC code [9, 7].
There are many different methods by which one can construct a quantum error correction code; however, for brevity and for the scope in this article, we will focus on stabilizer quantum codes, a class of codes for which the codespace $\mathsf{C}$ is defined by the simultaneous $+1$ eigenspace of Pauli stabilizer generators; that is, operators in the codespace, which take on the following form:
$\mathsf{C}=\{\ket{\psi}\in\mathcal{H}:S\ket{\psi}=\ket{\psi},\forall S\in\mathsf{S}\}~{}.$ (1)
The set of generators $\mathsf{S}=\{S\}$ which stabilize the codespace form an Abelian subgroup which is known as the stabilizer group $\mathsf{S}$.
In order to define the logical operators of a quantum code, we first must define several terms. The normalizer $\mathcal{N}$ of the stabilizer group is defined as
@@ II.1 Quantum Error Correction & Tensor Networks
Subsystem codes are typically stabilizer codes in which the encoded subspace is split up and subdivided. Normally, the Hilbert space of a stabilizer code is divided as $\mathcal{H}=\mathcal{H}_{C}\oplus\mathcal{H}_{\bar{C}}$, where $\mathcal{H}_{C}$ is used to store the logical information, and $\mathcal{H}_{\bar{C}}$ is the complement. We can additionally subdivide $\mathcal{H}_{C}$ further as $\mathcal{H}_{C}=\mathcal{H}_{L}\otimes\mathcal{H}_{J}$, where $\mathcal{H}_{L}$ is the logical subspace and $\mathcal{H}_{J}$ is known as the junk (or gauge) subspace [7]. The gauge subspace is used to help diagnose and correct errors inside of $\mathcal{H}_{L}$; such a feat is accomplished by defining the gauge group such that the stabilizer subgroup is the center $\mathsf{Z}(\mathsf{G})$ of the gauge group
$\mathsf{Z}(\mathsf{G}):=\{i\mathcal{I},\mathsf{S}\}=\mathsf{C}(\mathsf{G})\cap\mathsf{G}~{},$ (3)
where $\mathsf{G}$ is the junk (or gauge) group associated with $\mathcal{H}_{J}$, and $\mathsf{C}(\mathsf{G})$ is the centralizer of $\mathsf{G}$, which is defined as $\mathsf{C}(\mathsf{G}):=\{\mathcal{O}\in\mathsf{G}|[\mathcal{O},S]=0,\forall S\in\mathsf{S}\}$.
As a specific example of a stabilizer code, we introduce here the Steane code [56, 57], which is a Calderbank-Shor-Steane code. This $\llbracket 7,1,3\rrbracket$ code has X and Z stabilizers symmetrically split in the parity check matrix, and the stabilizer generators take the form
$X_{1}X_{4}X_{6}X_{7}~{},$ (4)
$X_{2}X_{4}X_{5}X_{7}~{},$ (5)
$X_{3}X_{5}X_{6}X_{7}~{},$ (6)
$Z_{1}Z_{4}Z_{6}Z_{7}~{},$ (7)
$Z_{2}Z_{4}Z_{5}Z_{7}~{},$ (8)
$Z_{3}Z_{5}Z_{6}Z_{7}~{},$ (9)
@@
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}\ ,$ (10)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}\ .$ (11)
Another small example of a quantum error correction code is the perfect $\llbracket 5,1,3\rrbracket$ code, which is known as the smallest $d=3$ code [58]; its stabilizer generators are
$X_{1}Z_{2}Z_{3}X_{4}~{},$ (12)
$X_{2}Z_{3}Z_{4}X_{5}~{},$ (13)
$X_{1}X_{3}Z_{4}Z_{5}~{},$ (14)
$Z_{1}X_{2}X_{4}Z_{5}~{},$ (15)
@@
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}~{},$ (16)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}~{}.$ (17)
These two examples in particular are important, as they constitute some of the elemental seed codes for constructing a holographic quantum code via the quantum LEGO (qLEGO) formalism [49, 59, 60]. More succinctly, the qLEGO formalism provides a method by which tensor networks can be employed in order to construct and simulate large QEC codes from seed codes. One starts by defining a quantum state as
|$\ket{\psi}=T_{i_{1}\cdots i_{m}}\ket{i_{1}\cdots i_{m}}~{},$ (18)
where the tensor $T_{i_{1}\cdots i_{m}}$ also can be used to describe a mapping from input subsystems to output subsystems, as shown in (a), (b) of Figure 1. Here the repeated indices are summed over. If qubits $1$ through $k$ are maximally entangled, we can write an isometry for $T_{i_{1}\cdots i_{m}}$ as
$V_{\ket{\bar{\psi}}}:=T_{i_{1}\cdots i_{m}}\ket{i_{k+1}\cdots i_{m}}\bra{i_{1}\cdots i_{k}}~{},$ (19)
in which we can easily take the incoming $i_{1}\cdots i_{k}$ input legs as the encoding qudits for an $\llbracket n,k\rrbracket$ quantum error correction code (where $n,k$ represent the numbers of physical and logically-encoded qudits, respectively), following [59]. In general, it is not necessary that $T_{i_{1}\cdots i_{m}}$ represent an isometry [59]; however, a convenient choice is to require that $T_{i_{1}\cdots i_{m}}$ corresponds to a multipartite maximally-entangled (MME) state [61], thereby necessitating that a many-body quantum state adhere to the condition
$\rho_{s}=\text{Tr}_{s^{c}}\big{[}\ket{\psi}\bra{\psi}\big{]}\propto\mathbb{I}~{},$ (20)
where $s\in\{1\cdots n\}$, $|s|\leq\left\lfloor{\frac{n}{2}}\right\rfloor$, and $s^{c}$ denotes the complementary set of subsystems to $s$. States adhering to this property are known as k-uniform, and have applications in many areas of quantum information theory [61]. Note that the qLEGO formalism also applies to more general tensors that do not form isometries, the contraction of isometric tensors (such as those above) permit the straightforward formation of a quantum code’s encoding circuit [59].
@@
Figure 1: Tensor-network diagrams depicting (a)-(b) the tensor $T_{i_{1}\cdots i_{m}}$ as it pertains to a quantum state, as well as an isometric map via the Choi-Jamiolkowski isomorphism [8], and (c)-(d) the action of tensor pushing an operator $\mathcal{O}$ through the input subsystems of the isometry in order to find an output-leg representation, which we name $\mathcal{O}^{\prime}$.
@@
Figure 2(a) shows the support of a unitary logical operator $\mathcal{\bar{O}}$ on the output subsystems as $\mathcal{\bar{O}}=\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$ (in our specific case, for simplicity). In (b), we see that contracting two such tensors along their respective indices gives rise to a new conjoined tensor (in this case with two logical qubits). Finding the new logical operators for this new isometry is quite simple, and follows the tensor-matching formalism described in detail in [59, 60]. As shown in (c), the principle requirement is that support on contracted output indices must be identical up to complex conjugation; in the specific case of Figure 2(c), the requirements must be that $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$.
@@
Figure 2: (a) One may define an operator $\mathcal{\bar{O}}$ on the input legs to a tensor by its representation (in the present case $\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$) on the output legs. (b) Contracting indices together yields a new, conjoined tensor network. (c) Deducing the new logical operator is possible by keeping the original logical operators that satisfy the matching condition $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$ [59, 60].
II.2 Holographic Quantum Codes

### v3

# and Decoding of Holographic Quantum Codes

@@ Abstract
Quantum error correction (QEC) is a crucial prerequisite for future large-scale quantum computation. Finding and analyzing new QEC codes, along with efficient decoding and fault-tolerance protocols, is central to this effort. Holographic codes are a recent class of generalized concatenated codes derived from holographic bulk/boundary dualities. In addition to exploring the physics of such dualities, these codes possess useful QEC properties such as tunable encoding rates, distance scaling competitive with other well-studied code classes, and excellent recovery thresholds. To allow for a comprehensive analysis of holographic code constructions, we introduce LEGO_HQEC, a software package utilizing the quantum LEGO formalism. This package allows for the construction and analysis of holographic codes on regular hyperbolic tilings, computing their stabilizer generators and logical operators for a specified number of seed codes and layers. Three decoders are included: an erasure decoder based on Gaussian elimination; an integer-optimization decoder; and a tensor-network decoder. With these tools, LEGO_HQEC enables systematic studies of both previously known holographic codes and novel variants. As a demonstration, we provide new numerical results on the holographic black-hole pentagon code, establishing its threshold behavior under the erasure channel as a benchmark example.
@@ I Introduction
The field of quantum computing is currently focused on understanding how best to construct a large-scale fault-tolerant quantum computer [campbell2017roadsterhal, gottesman2013fault, preskill2018quantum]. The physical levels of noise that dominate in current quantum computing hardware are simply too high to reliably run all but the simplest quantum algorithms. As a solution, quantum error correction (QEC) techniques [lidar2013quantum, wilde2013quantum, nielsen2010quantum, gottesman1997stabilizer] were proposed in order to orchestrate strategies that protect information at the logical level and facilitate the reliable execution of quantum algorithms on quantum devices.
@@
Although QEC in theory posits a method by which practically useful quantum computing could be realized, in reality the constraints imposed by current state-of-the-art code constructions and so-called fault-tolerance protocols are extremely resource-intensive [gottesmanFT]. As such, almost all areas of the QEC design pipeline are under active research. These research areas range from the theoretical to the practical, and include: code constructions [albert2022bosonic, breuckmann2021quantum, jahn_topical]; fault tolerant error correction protocols [between_shorsteane, dhruvi_groovy, chamberland2018flag, delfosse2021beyond, bombin2015single, divincenzo1996fault, divincenzo2007effective, chao2018quantum, aliferis2005quantum, knill2005quantum, shor1996fault, delfosse2020short]; fault-tolerant universal logical gate set implementations, such as magic-state distillation [bravyi2005universal], gauge-fixing [bombin2015gauge], and code switching [codeswitching1, codeswitching2, codeswitching3], among others [jochym2014using, chamberland2016thresholds]; decoder design [iolius2023decoding, dinur2023good, roffe2020decoding, babar2015fifteen]; architectural proposals [campbell2017roadsterhal, terhal2015quantum, yoshida2024concatenate]; among many other engineering-level challenges [fowler2012surface, breuckmann2021quantum].
@@
Holographic quantum codes are a recently introduced class of quantum error correction codes [preskill_happy_code, harlow2015bulk, jahn_topical], derived as models of bulk/boundary dualities in theoretical physics, in particular the AdS/CFT correspondences [maldacena1999large, witten1998anti]. The main idea for these codes lies in the fact that bulk degrees of freedom, which are associated with logical qubits in $d$ dimensions, are mapped to a (d1)(d-1)-dimensional boundary, with the hyperbolic structure of the bulk ensuring that the physical Hilbert space is larger than its logical counterpart. Constructed on hyperbolic manifold tessellations, they can be expressed as generalized concatenated codes and allow for extremely high (and tunable) encoding rates as well as excellent threshold performance in comparison to other leading quantum code constructions, such as quantum low-density parity-check (qLDPC) [breuckmann2021quantum] and topological (TQEC) quantum codes [fowler2012surface, terhal2015quantum]. However, for all practical purposes, actually generating the conformal quasicrystal lattices [conformal_quasicrystals_holography, jahn2022tensor_qcft, central_charge_aperiodic] responsible for defining physical-qubit representations of stabilizers and logical operators is a challenging task, exacerbated by the need to apply suitable decoding algorithms derived from tensor-network decoding techniques [joe_chubb, farrelly_parallel_decoding, farrelly_tn_codes, chubb2021statistical]. In this vein, having a software tool at hand to compute the effect of quantum noise channels is indispensable for the study of new holographic (as well as generalized concatenated) quantum codes.
In previous studies [farrelly_tn_codes, Farrelly_ltnc, qecsim], tensor-network decoders have been implemented using the Julia language (which is available at https://github.com/qecsim/TensorNetworkCodes.jl). Although one can in principle also build holographic tensor-network codes using this package, our work provides a specific advantage for studying holographic codes due to its streamlined setup and its specialization to generating regular hyperbolic lattices with a quasiperiodic boundary in an efficient manner, as well as its ability to test codes under a wide range of possible noise channels; these include the erasure, depolarizing, and biased-noise channels (such as those explored in [fan2024overcoming]).
@@
In this manuscript, we introduce a novel software tool named LEGO_HQEC in order to expedite and lower the barrier to entry for this new subfield of quantum error correction. In particular, we implement a stabilizer and logical operator generator that returns a list of these for both maximum-rate and asymptotically zero-rate variants of many existing holographic codes, as well as providing valuable customization options for constructing homogeneous or heterogeneous constructions (i.e., holographic codes constructed from either one or several smaller seed codes) on hyperbolic manifolds parameterized by Schläfli symbols $\{p,q{=}4\}$ [ratcliffe1994foundations]. In addition to this tool, we provide three different decoders which have been used in recent studies regarding the threshold properties of holographic codes for practical quantum computing: an erasure decoder built using a Gaussian elimination (GE) method [steinberg2024far, harris_css_hqec]; a Pauli-noise integer-optimization decoder, as from [harris_int_opt, steinberg2024far]; and a tensor-network decoder, which was utilized in [fan2024overcoming, farrelly_parallel_decoding, farrelly_tn_codes] additionally for Pauli noise. With the aim of developing newer and better open-source software tools for studying quantum error correction, this package has been made publicly available at https://github.com/QML-Group/HQEC. In this way, the overarching design premise of LEGO_HQEC involves automation and extendability for creating, analyzing, and decoding holographic quantum codes.
@@
This paper is organized as follows: Section II gives an overview of the quantum LEGO formalism, and how it relates to stabilizer and holographic QEC codes; Section III details the structure of LEGO_HQEC, starting with boundary stabilizer and logical operator generation, and ending with a concise overview of the distinct decoders offered; in Section IV, we show two examples using: the $\llbracket 5,1,3\rrbracket$ pentagon (or Harlow-Preskill-Pastawski-Yoshida (HaPPY) [preskill_happy_code]) code, as well as a holographic Steane code [harris_css_hqec]. We further present novel numerical results on the black-hole HaPPY code in Section IV-A1, highlighting its threshold behavior under the erasure channel as a demonstration of the tool’s utility. The calculated runtime for generating stabilizers and logical operators for several layers of the asymptotically zero-rate HaPPY code is presented, as well as decoding for erasure and Pauli noise in the asymptotically zero-rate Steane code, in addition to the maximum-rate HaPPY code. We conclude by proposing several extensions to LEGO_HQEC in Section VI.
@@ II Background
II-A Quantum Error Correction & Tensor Networks
Quantum error correction (QEC) codes are used to protect information at the level of physical qubits (or qudits) from deleterious environmental noise effects. Such protection is accomplished by redundantly mapping physical qubit states into subspaces of a larger, multipartite Hilbert space; this description at the level of an encoding isometry effectively defines a QEC code [nielsen2010quantum, lidar2013quantum].
There are many different methods by which one can construct a QEC code; however, for brevity and for the scope in this article, we will focus on quantum stabilizer codes, a class of codes for which the codespace $\mathsf{C}$ is defined by the simultaneous +1+1 eigenspace of Pauli stabilizer generators; that is,
[…]
wherein the set of generators $\{S\}$ which stabilize the codespace form an Abelian subgroup which is known as the stabilizer group $\mathsf{S}$, and as such $\{S\}\in\mathsf{S}$. In order to define the logical operators of a quantum code, we first must define several terms. The normalizer $\mathcal{N}$ of the stabilizer group is defined as
@@ II-A Quantum Error Correction & Tensor Networks
As a specific example of a stabilizer code, we introduce here the Steane code [steanecode1, steanecode2], which belongs to the class of Calderbank-Shor-Steane codes. This $\llbracket 7,1,3\rrbracket$ code has $X$ and $Z$ stabilizers symmetrically split in the parity check matrix, and the stabilizer generators take the form
$X_{1}X_{4}X_{6}X_{7}~,$ (3a)
$X_{2}X_{4}X_{5}X_{7}~,$ (3b)
$X_{3}X_{5}X_{6}X_{7}~,$ (3c)
$Z_{1}Z_{4}Z_{6}Z_{7}~,$ (3d)
$Z_{2}Z_{4}Z_{5}Z_{7}~,$ (3e)
$Z_{3}Z_{5}Z_{6}Z_{7}~,$ (3f)
@@
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}\ ,$ (4a)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}\ .$ (4b)
Another small example of a QEC code is the perfect $\llbracket 5,1,3\rrbracket$ code, which is known as the smallest code with code distance $d=3$ [laflamme1996perfect]; its stabilizer generators are
$X_{1}Z_{2}Z_{3}X_{4}~,$ (5a)
$X_{2}Z_{3}Z_{4}X_{5}~,$ (5b)
$X_{1}X_{3}Z_{4}Z_{5}~,$ (5c)
$Z_{1}X_{2}X_{4}Z_{5}~,$ (5d)
@@
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}~,$ (6a)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}~.$ (6b)
These two examples in particular are important, as they constitute some of the elemental seed codes for constructing a holographic quantum code via the quantum LEGO (qLEGO) formalism [farrelly_tn_codes, quantumlego, Farrelly_ltnc]. More succinctly, the qLEGO formalism provides a method by which tensor networks can be employed in order to construct and simulate large QEC codes from seed codes. One starts by defining a quantum state as
[…]
where the tensor $T_{i_{1}\cdots i_{m}}$ also can be used to describe a mapping from input subsystems to output subsystems, as shown in Figure 1(a)-(b). Here the repeated indices are summed over. If qubits $1$ through $k$ are maximally entangled, we can write an isometry for $T_{i_{1}\cdots i_{m}}$ as
[…]
in which we can easily take the incoming $i_{1}\cdots i_{k}$ input legs as the encoding qudits for an $\llbracket n,k\rrbracket$ QEC code (where $n$ and $k$ represent the number of physical and logically-encoded qudits, respectively), following [quantumlego]. In general, it is not necessary that $T_{i_{1}\cdots i_{m}}$ represent an isometry [quantumlego]; however, a convenient choice is to require that $T_{i_{1}\cdots i_{m}}$ corresponds to a multipartite maximally-entangled (MME) state [zyczkowski_mme], thereby necessitating that a many-body quantum state adheres to the condition
[…]
where $s\in\{1\cdots n\}$, $|s|\leq\left\lfloor{\frac{n}{2}}\right\rfloor$, and $s^{c}$ denotes the subsystems set complementary to $s$. States adhering to this property are known as k-uniform, and have applications in many areas of quantum information theory [zyczkowski_mme]. Note that the qLEGO formalism also applies to more general tensors that do not form isometries, as the contraction of isometric tensors (such as those above) permits the straightforward formation of a quantum code’s encoding circuit [quantumlego].
@@
Figure 1: Tensor-network diagrams depicting (a)-(b) the tensor $T_{i_{1}\cdots i_{m}}$ as it pertains to a quantum state, as well as an isometric map via the Choi-Jamiolkowski isomorphism [wilde2013quantum]. Here, red open legs represent logical input qubits, and black legs depict physical output qubits. (c)-(d) The action of tensor pushing an operator $\mathcal{O}$ through the input subsystems of the isometry in order to find an output-leg representation, which we name $\mathcal{O}^{\prime}$.
@@
Figure 2(a) shows the support of a unitary logical operator $\mathcal{\bar{O}}$ on the output subsystems as $\mathcal{\bar{O}}=\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$ (in our specific case, for simplicity). In (b), we see that contracting two such tensors along their respective indices gives rise to a new conjoined tensor (in this case with two logical qubits). Finding the new logical operators for this new isometry is quite simple, and follows the tensor-matching formalism described in detail in [quantumlego, Farrelly_ltnc]. As shown in (c), the principle requirement is that support on contracted output indices must be identical up to complex conjugation; in the specific case of Figure 2(c), the requirements must be that $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$.
@@
Figure 2: (a) One may define an operator $\mathcal{\bar{O}}$ on an input leg to a tensor by its representation (in the present case $\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$) on the output legs. (b) Contracting indices together yields a new, conjoined tensor network. (c) Deducing the new logical operator is possible by keeping the original logical operators that satisfy the matching condition $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$ [quantumlego, Farrelly_ltnc].
The concepts and operations explained above will be of paramount use in the software implementation, as we will see in Section III.
II-B Holographic Quantum Codes